A laser beam is incident at an angle of 33.0° to the vertical onto a solution of cornsyrup in water.
(a) If the beam is refracted to 24.84° to the vertical, what is the index ofrefraction of the syrup solution?
1

(b) Suppose the light is red, with wavelength 632.8 nm in a vacuum.Find its wavelength in the solution.
2 nm

(c) What is its frequencyin the solution?
3 Hz

(d) What is its speed in the solution?
4 m/s
Answer #1
Concepts and reason

The concepts used to solve this problem are Snell’s law, index of refraction, Planck–Einstein relation, and frequency of light.

First, use the expression for Snell’s law to determine the relation between the refractive index of a medium and the angle of incidence and refraction.

Use the relation between the wavelength and refractive index at the interface to determine the wavelength of the laser in corn syrup.

Use the Planck–Einstein relation to determine the relation between the frequency of laser and its wavelength.

Finally, use the relation between the refractive index and the speed of light to determine its speed in corn syrup solution.

Fundamentals

According to Snell’s law, the ratio of the sine of the angle of incidence and the sine of the angle of refraction is constant.

The expression for the Snell’s law is given below:

n1n2=sinθ2sinθ1\frac{{{n_1}}}{{{n_2}}} = \frac{{\sin {\theta _2}}}{{\sin {\theta _1}}}

Here, the refractive index of vacuum is n1{n_1} , the refractive index of the second medium is n2{n_2} , the incident angle is θ1{\theta _1} , and the refracted angle is θ2{\theta _2} .

The expression for the relation between the refractive indices and the wavelength of light in a medium is given below:

n1n2=λ2λ1\frac{{{n_1}}}{{{n_2}}} = \frac{{{\lambda _2}}}{{{\lambda _1}}}

Here, the wavelength of light in the vacuum is λ1{\lambda _1} , and the wavelength of light in the second medium is λ2{\lambda _2} .

The expression for the Planck-Einstein relation is given below:

f=cλ1f = \frac{c}{{{\lambda _1}}}

Here, the frequency of light is ff , and the speed of light in vacuum is cc .

The expression relating to the speed of the light and the refractive index of a medium is given below:

vc=n1n2\frac{v}{c} = \frac{{{n_1}}}{{{n_2}}}

Here, the speed of laser in the second medium is vv , the speed of light in vacuum is cc , the refractive index of the vacuum is n1{n_1} , the refractive index of the second medium is n2{n_2} .

(a)

The expression for Snell’s law in terms of refractive index, incident angle and refracted angle is given below:

n1n2=sinθ2sinθ1\frac{{{n_1}}}{{{n_2}}} = \frac{{\sin {\theta _2}}}{{\sin {\theta _1}}}

Here, n2{n_2} is the refractive index of the corn syrup.

Rearrange the above expression for the refractive index of the corn syrup,

n2=n1(sinθ1sinθ2){n_2} = {n_1}\left( {\frac{{\sin {\theta _1}}}{{\sin {\theta _2}}}} \right)

Substitute 11 for n1{n_1} , 3333^\circ for θ1{\theta _1} , 24.8424.84^\circ for θ2{\theta _2}

n2=(1)(sin(33)sin(24.84))=1.296\begin{array}{c}\\{n_2} = \left( 1 \right)\left( {\frac{{\sin \left( {33} \right)}}{{\sin \left( {24.84} \right)}}} \right)\\\\ = 1.296\\\end{array}

[Part a

Part a

[Part a]

The relation between the wavelength in a medium and its refractive index is given below:

λα1n\lambda \alpha \frac{1}{n}

The ratio between the refractive indices between the vacuum and corn syrup is given below:

n1n2=λ2λ1\frac{{{n_1}}}{{{n_2}}} = \frac{{{\lambda _2}}}{{{\lambda _1}}}

Here, λ2{\lambda _2} is the wavelength of the light in the corn syrup.

Rearrange the above expression for the wavelength in the corn syrup,

λ2=λ1(n1n2){\lambda _2} = {\lambda _1}\left( {\frac{{{n_1}}}{{{n_2}}}} \right)

Substitute 11 for n1{n_1} , 1.2961.296 for n2{n_2} , and 632.8×109m632.8 \times {10^{ - 9}}\,{\rm{m}} for λ1{\lambda _1}

λ2=(632.8×109m)(11.296)=488.3×109m=488.3nm\begin{array}{c}\\{\lambda _2} = \left( {632.8 \times {{10}^{ - 9}}\,{\rm{m}}} \right)\left( {\frac{1}{{1.296}}} \right)\\\\ = 488.3 \times {10^{ - 9}}\,{\rm{m}}\\\\ = {\rm{488}}{\rm{.3}}\,{\rm{nm}}\\\end{array}

(c)

The expression for Planck-Einstein relation is given below:

f=cλ1f = \frac{c}{{{\lambda _1}}}

Substitute 3×108m/s3 \times {10^8}\,{\rm{m/s}} for cc , and 632.8×109m632.8 \times {10^{ - 9}}\,{\rm{m}} for λ1{\lambda _1}

f=3×108m/s632.8×109m=4.74×1014Hz\begin{array}{c}\\f = \frac{{3 \times {{10}^8}\,{\rm{m/s}}}}{{632.8 \times {{10}^{ - 9}}\,{\rm{m}}}}\\\\ = 4.74 \times {10^{14}}\,{\rm{Hz}}\\\end{array}

(d)

The expression relating to the refractive index and the speed of light is given below:

n1n2=v2v1\frac{{{n_1}}}{{{n_2}}} = \frac{{{v_2}}}{{{v_1}}}

Rearrange the above expression for the speed of light in the solution as shown below:

v2=v1(n1n2){v_2} = {v_1}\left( {\frac{{{n_1}}}{{{n_2}}}} \right)

Substitute 3×108ms13 \times {10^8}\,{\rm{m}}{{\rm{s}}^{ - 1}} for v1{v_1} , 11 for n1{n_1} , and 1.2961.296 for n2{n_2}

v2=(3×108ms1)(11.296)=2.314×108ms1\begin{array}{c}\\{v_2} = \left( {3 \times {{10}^8}\,{\rm{m}}{{\rm{s}}^{ - 1}}} \right)\left( {\frac{1}{{1.296}}} \right)\\\\ = 2.314 \times {10^8}\,{\rm{m}}{{\rm{s}}^{ - 1}}\\\end{array}

Ans: Part a

The refractive index of the corn syrup is 1.296{\bf{1}}{\bf{.296}} .

Part b

The wavelength of the red light in the corn syrup is 488.3nm{\bf{488}}{\bf{.3}}\,{\bf{nm}} .

Part c

The frequency of the light in the corn solution is 4.74×1014Hz{\bf{4}}{\bf{.74 \times 1}}{{\bf{0}}^{{\bf{14}}}}\,{\bf{Hz}} .

Part d

The speed of light in the corn syrup is 2.314×108ms1{\bf{2}}{\bf{.314 \times 1}}{{\bf{0}}^{\bf{8}}}\,{\bf{m}}{{\bf{s}}^{{\bf{ - 1}}}} .